OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (7, 0) and a directrix at x = -7

OpenStudy (anonymous):

Is it: y = 1/28 x^2 ?

random231 (random231):

nah!

OpenStudy (anonymous):

No? Oh. -28y=x^2? @random231

random231 (random231):

no you got the axis wrong.

random231 (random231):

it shud be y^2=28x

OpenStudy (anonymous):

That's not an option though. These are the options: a. y 1/28 = x^2 b. x 1/28 = y^2 c. -28y = x^2

OpenStudy (anonymous):

@random231

OpenStudy (cwrw238):

standard form can be written y^2 = 4 a x where the focus is at ( a,0) that would make it y" = 28x but i'd have to check that out

OpenStudy (anonymous):

But how am I supposed to answer the question?

OpenStudy (cwrw238):

if the directrix is x = -7 and focus is at (7,0) that means the parabola opens to the right and has the general equation y^2 = 4 a x where (a,0) is the focus here a = 7 so the rquired equation is y^2 = 4*7x ie y^2 = 28x

OpenStudy (cwrw238):